A party host gives a door prize to one guest chosen at random. There are 47 men and 40 women at the party. What is the probability that the prize goes to a woman?
1. glenbenn50@yahoo.com

Answer by jim_thompson5910(28536)

Since there are 47 men and 40 women, this means that there are 47+40=87 people total.

Probability prize goes to woman = # of women/# total = 40/87 = 0.4597

which is about a 45.97% chance